# Rational Gluing in Edge Replacement Systems

### v1

@@ Abstract.
In this paper, we prove the rationality of the gluing relation of edge replacement systems, which were introduced for studying rearrangement groups of fractals. More precisely, we describe an algorithmic procedure for building a finite state automaton that recognizes pairs or equivalent sequences that are glued in the fractal. This fits in recent interest towards the rationality of gluing relations on totally disconnected compact metrizable spaces.
@@ Introduction
In this paper we study the limit space of edge replacement systems, which arose from the context of rearrangement groups of fractals introduced in [BF19]. This is a family of Thompson-like groups acting on fractals which have now been studied and used in several papers [WZ19, Tar24, PT22, Ner23, Tar23, BF23, Tar23a]. We will construct automata that recognize pairs of infinite words precisely when they are equivalent under the gluing relation that defines the limit space, meaning that the gluing relation is rational. Different ways to define rationality have been employed in the literature, especially in symbolic dynamics with notions like edge and sofic shifts. By the well-known Alexandroff-Hausdorff Theorem, every compact metrizable space $X$ is a quotient of the Cantor space $\mathfrak{C}$. When seeing $\mathfrak{C}$ as a set of infinite words on a finite alphabet, this defines a coding on $X$: every point of $X$ is associated to an equivalence class of infinite words of $\mathfrak{C}$ that are “glued” by the equivalence relation. More generally, the Cantor space $\mathfrak{C}$ can replaced by a one-sided edge shift, which is the set of infinite words that can be read as walks on a finite directed graph. A natural and simple example of a coding is the binary coding of the unit interval, where each number in $[0,1]$ has one or two codings depending whether it is irrational or rational. While a general coding can be quite unwieldy, interesting and useful codings generally arise when there is an explicit rule for gluing two infinite words together. What motivates the interest in the rationality of a gluing relation is thus that there is indeed a rule that described when two words glue together, and that rule is “computable”.
@@
There are several examples of this in literature, tough they may be addressed with different names or be stronger or weaker versions of what we are describing here. The Gromov boundary of a hyperbolic group has rational gluing with respect to different codings. In [CP93, Paw15], it is proved that it is semi-Markovian with respect to the boundary of the tree of geodesics and other trees introduced by Gromov himself ([Gro87]). Being semi-Markovian is generally a stronger notion than rationality (see again [Paw15]). In [Per23], it is shown that the gluing associated to the horofunction boundary is rational. Limit spaces of contracting self similar groups have rational gluing with respect to the regular tree on which the groups act on (see [Nek07]). In [Ban23], Bandt explores the opposite point of view, building self-similar spaces starting from automata that define gluing relations and our construction fits in Bandt’s formalism.
@@
Recently in [BF19], for their study in group theory, Belk and Forrest introduced edge replacement systems, which are graph rewriting systems that produce a plethora of different self-similar fractals, such as the basilica [BF15] and airplane Julia sets [Tar24] and Ważewski dendrites [Tar23a]. As we will recall in the background Section, these fractals are obtained as quotients of the Cantor space under an equivalence “gluing” relation. Here we prove that this relation is rational, as conjectured in [Per23, Example 7.14]. We provide an explicit construction that associates to any edge replacement system a finite state automaton that recognizes pairs of equivalent infinite words.
@@ 1. Background on Edge Replacement Systems
The definitions given in this Section are the same of [BF19], except that we associate distinguished colors to loops in Definition 1.1. For a complete treatment of this topics we refer to [BF19].
@@ 1.1. Edge Replacement Systems
When the set of colors is a singleton, we will denote by $1$ the only color and we will essentially omit $\mathrm{Col}$, $\mathrm{c}$ and any related index from the notation.
@@
  • each replacement graph $\Gamma_{\mathrm{c}}$ is equipped with one or two boundary vertices: either two distinct vertices $\iota_{\mathrm{c}}$ and $\tau_{\mathrm{c}}$ or a single vertex $\lambda_{\mathrm{c}}$.
We require the following conditions:
  1. (A)
    if $\Gamma_{\mathrm{c}}$ has a single boundary vertex $\lambda_{\mathrm{c}}$, then each $\mathrm{c}$-colored edge among the base and replacement graphs is a loop;
  2. (B)
    if $\Gamma_{\mathrm{c}}$ has two boundary vertices $\iota_{\mathrm{c}}$ and $\tau_{\mathrm{c}}$, then each $\mathrm{c}$-colored edge among the base and replacement graphs is not a loop.
Examples of replacement systems are depicted in Figures 1, LABEL:, 3, LABEL: and 2.
{tikzpicture}
@@
{tikzpicture}
Figure 3. The Basilica replacement system $\mathcal{B}$.
A replacement system defines a rewriting system in the following way. We can expand the base graph $\Gamma_{0}$ by replacing one of its edges $e$ with the $\mathrm{c}(e)$-colored replacement graph $\Gamma_{\mathrm{c}}$, attaching the boundary vertices of $\Gamma_{\mathrm{c}}$ in place of the initial and terminal vertices $\iota_{\mathrm{c}}(e)$ and $\tau_{\mathrm{c}}(e)$ of the edge $e$. Note that, because of requirements (A) and (B) of Definition 1.1, this is coherent with the amount of vertices that need to be attached (one for loops, two otherwise). This very same procedure can be applied to any graph expansion of the base graph, which allows to obtain further graph expansions. For example, Figure 4 depicts graph expansions of the replacement systems $\mathcal{I}$ and $\mathcal{B}$ from Figures 3, LABEL: and 1, respectively.
{tikzpicture}
@@ 1.2.2. The Limit Space
Two sequences $\alpha$ and $\beta$ of $\Omega_{\mathcal{R}}$ are equivalent under the gluing relation if $\alpha_{1}\dots\alpha_{n}$ and $\beta\dots\beta_{n}$ are incident on a common vertex for every large enough $n\in\mathbb{N}$. In this case, we write $\alpha\sim\beta$.
This relation $\sim$ is supposed to glue together those sequences that represent the same point in the limit. If two distinct sequences $\alpha$ and $\beta$ of $\Omega_{\mathcal{R}}$ are glued by $\sim$, then they must share a (possibly trivial) common prefix $x$, after which $x_{1}\dots x_{k}\alpha_{k+1}\dots\alpha_{m}$ and $x_{1}\dots x_{k}\beta_{k+1}\dots\beta_{m}$ must be incident on a common vertex for all $m\geq k$. For example, in the interval replacement system $\mathcal{I}$, for every finite prefix $p$ one has $p0\bar{1}\sim p1\bar{0}$, and these are all the gluings on $\mathcal{I}$.
@@
    there is no edge between the boundary vertices of any replacement graph;
@@
    each replacement graph has at least two edges and at least one vertex that is not a boundary vertex.
@@
By [BF19, Theorem 1.25], limit spaces are non-empty compact metrizable spaces. The limit space for the replacement system $\mathcal{I}$ from Figure 1 is the (topological) interval. The limit space of $\mathcal{D}_{n}$ from Figure 2 is the Ważewski dendrite $D_{n}$ of order $n$, and Figure 4(a)111image in Figure 4(a) by Adam Majewski, licensed under CC BY 3.0 depicts $D_{3}$; see [Tar23a] for more details on these spaces and their rearrangement groups. The limit space of $\mathcal{B}$ from Figure 3 is the Basilica limit space, which is homeomorphic to the Julia set for the quadratic polynomial map $z^{2}-1$, depicted in Figure 4(b)222image in Figure 4(b) by Jim Belk and Bradley Forrest from [BF15]; see [BF15, BF19] for more details on the limit space and its rearrangement group.
@@ 2. The Gluing Automaton
Let $\Omega$ be a set of infinite sequences. We say that an equivalence relation on $\Omega$ is rational when there exists an automaton that recognizes a pair $(w,v)$ is and only if the two elements $w,v\in\Omega$ are related.
@@
Now consider an expanding replacement system $\mathcal{R}$ with colors $\mathrm{Col}=\{1,\dots,k\}$, base graph $\Gamma_{0}$ and replacement graphs $\Gamma_{1},\dots,\Gamma_{k}$. The gluing automaton $\mathrm{Gl}_{\mathcal{R}}=(\Sigma^{2},Q,\rightarrow,q_{0}(0))$ is built as follows.
@@
When processing one of the two entries of the $l$-th letter of a word in this alphabet, it will be natural to refer to that entry as if it were an edge of the $l$-th graph of the full expansion sequence (i.e., the sequence of graphs obtained by expanding, at each step, every edge of the previous graph, starting with the base graph). In this way, the two entries of the $l$-th letter of a word in $\Sigma^{2}$ represent two edges of the same graph, so in what follows we will simply describe their possible adjacency in the $l$-th graph of the full expansion sequence without explicitly mentioning the graph itself.
@@
The following Subsection will describe in more details what these states mean. The intuition is that, while processing the edges of our two sequences, the first set $Q_{0}$ consists of an initial state $q_{0}(0)$ along with a state $q_{0}(i)$ for each color $i$ that indicates that the two entries being processed are the same $i$-colored edge, while the second set $Q_{1}$ has a state for each quadruple $\big{(}\begin{smallmatrix}i&\gamma\\ j&\delta\end{smallmatrix}\big{)}$ where $i,j$ are the colors of the edges and $\gamma,\delta$ distinguish between the types of adjacency between the edges. The symbols $\mathrm{lp}$ and $\mathrm{db}$ stand for loop and double, respectively, and their meaning will be made clear when describing the transition functions.
@@
$q_{0}(i)\xrightarrow{(a,b)}q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\alpha\\ \mathrm{col}(b)&\beta\end{smallmatrix}\big{)}$ for all $i\in\{0\}\cup\mathrm{Col}$ and $a\neq b$ in $\Gamma_{i}^{1}$, where $\alpha$ and $\beta$ depend on the adjacency between $a$ and $b$ in $\Gamma_{i}$ as shown in Table 1.
@@
$q_{1}\big{(}\begin{smallmatrix}i&\gamma\\ j&\delta\end{smallmatrix}\big{)}\xrightarrow{(a,b)}q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\alpha\\ \mathrm{col}(b)&\beta\end{smallmatrix}\big{)}$ for all $i,j\in\{0\}\cup\mathrm{Col}$ whenever $a\in\Gamma_{i}^{1}$ and $b\in\Gamma_{j}^{1}$ satisfy the conditions described in Table 2, which also depend of $\gamma$ and $\delta$ and determine $\alpha$ and $\beta$.
The meaning of these transitions (especially Type 2) will be made clear in Section 3. The idea is that Type 0 transitions simply induce a copy of the color graph (Definition 1.2), Type 1 transitions arise when two sequences first diverge and Type 2 transitions describe the adjacency of sequences after their divergence.
Note that, when processing the pair $(a,b)$ in a state $q_{0}(i)$, there is a Type 1 transition if and only if $a$ and $b$ are adjacent in $\Gamma_{i}$. Moreover, in Type 2 transitions $\alpha$ and $\beta$ will never be $\mathrm{db}^{+}$ nor $\mathrm{db}^{-}$, since the replacement system is expanding.
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{in}\\ \mathrm{col}(b)&\mathrm{out}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{in}\\ \mathrm{col}(b)&\mathrm{in}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{out}\\ \mathrm{col}(b)&\mathrm{out}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{out}\\ \mathrm{col}(b)&\mathrm{in}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{in}\\ \mathrm{col}(b)&\mathrm{lp}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{out}\\ \mathrm{col}(b)&\mathrm{lp}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{lp}\\ \mathrm{col}(b)&\mathrm{out}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{lp}\\ \mathrm{col}(b)&\mathrm{in}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{lp}\\ \mathrm{col}(b)&\mathrm{lp}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{db}^{+}\\ \mathrm{col}(b)&\mathrm{db}^{+}\end{smallmatrix}\big{)}$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{db}^{+}\\ \mathrm{col}(b)&\mathrm{db}^{-}\end{smallmatrix}\big{)}$
Table 1. Type 1 transitions and the types of adjacency.
@@
Table 2. Type 2 transitions. Each entry $\begin{smallmatrix}\alpha\\ \beta\end{smallmatrix}$ represents a transition to the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\alpha\\ \mathrm{col}(b)&\beta\end{smallmatrix}\big{)}$. An empty entry at $\big{(}\begin{smallmatrix}\gamma\\ \delta\end{smallmatrix}\big{)}$ means that the state $q_{1}\big{(}\begin{smallmatrix}i&\gamma\\ j&\delta\end{smallmatrix}\big{)}$ does not exist.
For example, Figure 6 depicts the gluing automaton $\mathrm{Gl}_{\mathcal{I}}$ for the interval replacement system from Figure 1 (which is a classical example) and Figure 7 depicts the gluing automaton $\mathrm{Gl}_{\mathcal{D}_{n}}$ for the Dendrite replacement systems from Figure 2. Finally, Table 3 represents the gluing automaton of the Basilica replacement system: the automaton $\mathrm{Gl}_{\mathcal{B}}$ features a transition $q^{(1)}\xrightarrow{(a,b)}q^{(2)}$ if and only if Table 3 features a row that reads $q^{(1)}\mid(a,b)\mid q^{(2)}$.
For example, to deduce the transition $q_{1}\big{(}\begin{smallmatrix}1&\mathrm{in}\\ 1&\mathrm{out}\end{smallmatrix}\big{)}\xrightarrow{(1,0)}q_{1}\big{(}\begin{smallmatrix}1&\mathrm{in}\\ 1&\mathrm{out}\end{smallmatrix}\big{)}$ in Figure 6 from Table 2 one needs to observe that $\gamma=\mathrm{in}$ and $\tau(1)$ is the terminal vertex $\tau_{1}$ of the replacement graph, so one needs to consider the second column, and that $\delta=\mathrm{out}$ and $\iota(0)$ is the initial vertex $\iota_{1}$ of the replacement graph, so one needs to consider the fourth row. This entry of the table reads $\begin{smallmatrix}\mathrm{in}\\ \mathrm{out}\end{smallmatrix}$.
{tikzpicture}
@@
{tikzpicture}

### v2

@@ Abstract.
In this paper, we prove the rationality of the gluing relation of edge replacement systems, which were introduced for studying rearrangement groups of fractals. More precisely, we describe an algorithmic procedure for building a finite state automaton that recognizes pairs of equivalent sequences that are glued in the fractal. This fits in recent interest towards the rationality of gluing relations on totally disconnected compact metrizable spaces.
@@ Introduction
In this paper we study the limit space of edge replacement systems, which arose from the context of rearrangement groups of fractals introduced in [BF19]. This is a family of Thompson-like groups acting on fractals which have now been studied and used in several papers [MR4033504, airplane, IG, BasilicaDense, RearrConj, Quasisymmetries, dendrite]. We will construct automata that recognize pairs of infinite words precisely when they are equivalent under the gluing relation that defines the limit space, meaning that the gluing relation is rational. Different ways to define rationality have been employed in the literature, especially in symbolic dynamics with notions like edge and sofic shifts. By the well-known Alexandroff-Hausdorff Theorem, every compact metrizable space $X$ is a quotient of the Cantor space $\mathfrak{C}$. When seeing $\mathfrak{C}$ as a set of infinite words on a finite alphabet, this defines a coding on $X$: every point of $X$ is associated to an equivalence class of infinite words of $\mathfrak{C}$ that are “glued” by the equivalence relation. More generally, the Cantor space $\mathfrak{C}$ can replaced by a one-sided edge shift, which is the set of infinite words that can be read as walks on a finite directed graph. A natural and simple example of a coding is the binary coding of the unit interval, where each number in $[0,1]$ has either one or two binary codings depending on whether it is dyadic. While a general coding can be quite unwieldy, interesting and useful codings generally arise when there is an explicit rule for gluing two infinite words together. What motivates the interest in the rationality of a gluing relation is thus that there is indeed a rule that describes when two words glue together, and that rule is “computable”.
@@
There are several examples of this in literature, though they may be addressed with different names or be stronger or weaker versions of what we are describing here. The Gromov boundary of a hyperbolic group has rational gluing with respect to different codings. In [CP, P], it is proved that it is semi-Markovian with respect to the boundary of the tree of geodesics and other trees introduced by Gromov himself ([G]). Being semi-Markovian is generally a stronger notion than rationality (see again [P]). In [Perego], it is shown that the gluing associated to the horofunction boundary is rational. Limit spaces of contracting self similar groups have rational gluing with respect to the regular tree on which the groups act (see [N2]). In [Bandt], Bandt explores the opposite point of view, building self-similar spaces starting from automata that define gluing relations and our construction fits in Bandt’s formalism.
@@
Recently in [BF19], for their study in group theory, Belk and Forrest introduced edge replacement systems, which are graph rewriting systems that produce a plethora of different self-similar fractals, such as the basilica [BF15] and airplane Julia sets [airplane] and Ważewski dendrites [dendrite]. As we will recall in the background Section, these fractals are obtained as quotients of the Cantor space under an equivalence “gluing” relation. Here we prove that this relation is rational, as conjectured in [Perego, Example 7.14]. We provide an explicit construction that associates to any edge replacement system a finite state automaton that recognizes pairs of equivalent infinite words.
@@ 1. Background on Edge Replacement Systems
The definitions given in this Section are the same of [BF19], except that we will associate distinguished colors to loops as explained in Subsection 1.1.1 (this change is only technical and does not affect the limit spaces that can be achieved). For a complete treatment of this topic we refer to [BF19].
@@ 1.1. Edge Replacement Systems
When the set of colors is a singleton, we will denote by $1$ the only color and we will essentially omit $\mathrm{Col}$, $\mathrm{c}$ and any related index from the notation. Moreover, given a graph $\Gamma$ it will often be useful to denote the set of edges of $\Gamma$ by $\Gamma^{1}$.
@@
  • each replacement graph $\Gamma_{\mathrm{c}}$ is equipped with two distinct boundary vertices $\iota_{\mathrm{c}}$ and $\tau_{\mathrm{c}}$.
Examples of replacement systems are depicted in Figures 1, LABEL: and 2.
$\Gamma_{0}=$$S$$\Gamma_{1}=$$\iota$$\tau$001111
@@
1.1.1. Colors and Loops
Throughout this paper, we will make the following assumption on the graph expansions of our edge replacement systems, as it will make it easier to build our gluing automaton.
Assumption 1.2.
We assume that edges of the same color are either all loops or all not loops.
Under this assumption, it is natural to identify the two boundary vertices for colors that represent loops. Each replacement graph $\Gamma_{c}$ will then equipped with either one or two boundary vertices: a single vertex $\lambda_{c}$ for loops and two distinct vertices $\iota_{c}$ and $\tau_{c}$ for non-loops.
Remark 1.3.
This assumption is not really a restriction on the limit spaces that we can produce, nor on their rearrangement groups. Indeed, given a generic edge replacement system, we can always split a color $c$ into two colors $c_{\text{loop}}$ and $c_{\text{non-loop}}$ and create a new edge replacement system that produces the same limit space and the same rearrangement group. This creates a replacement system that satisfies 1.2 because whether an edge is a loop or not only depends on whether it is a loop of the base graph or of some replacement graph. For example, the original basilica replacement system Figure 3 becomes the one portrayed in Figure 4.
More precisely, the new edge replacement system $\mathcal{R}^{*}$ is obtained by replacing each color $c$ by two new colors $c_{\text{loop}}$ and $c_{\text{non-loop}}$, where $c_{\text{loop}}$ is used on loops that were formerly of color $c$ and $c_{\text{non-loop}}$ is used on every other edge of color $c$. Even if we will not discuss rearrangement groups here, it is also not hard to see that this modification does not affect the rearrangement group.
$\Gamma_{0}=$$L$$R$$\Gamma_{1}=$$\iota_{1}$$\tau_{1}$111122223333
Figure 3. The original Basilica replacement system.
[…]
Figure 4. The modified Basilica replacement system $\mathcal{B}$.
1.1.2. Graph Expansions
A replacement system defines a rewriting system in the following way. We can expand the base graph $\Gamma_{0}$ by replacing one of its edges $e$ with the $\mathrm{c}(e)$-colored replacement graph $\Gamma_{\mathrm{c}(e)}$, attaching the boundary vertices of $\Gamma_{\mathrm{c}(e)}$ in place of the initial and terminal vertices $\iota_{\mathrm{c}}(e)$ and $\tau_{\mathrm{c}}(e)$ of the edge $e$. Note that, because we identify the boundary vertices for colors associated to loops as explained in Subsection 1.1.1, this is coherent with the amount of vertices that need to be attached (one for loops, two otherwise). This very same procedure can be applied to any graph expansion of the base graph, which allows us to obtain further graph expansions. For example, Figure 5 depicts graph expansions of the replacement systems $\mathcal{I}$ and $\mathcal{B}$ from Figures 4, LABEL: and 1, respectively. This prompts the following useful definition.
Definition 1.4.
The full expansion sequence is the sequence of graphs $E_{m}$ obtained by expanding, at each step, every edge of the previous graph, starting with the base graph $E_{0}=\Gamma_{0}$.
0010010010010010110110110111111111
@@ 1.2.2. The Limit Space
Two sequences $\alpha$ and $\beta$ of $\Omega_{\mathcal{R}}$ are equivalent under the gluing relation if $\alpha_{1}\dots\alpha_{n}$ and $\beta_{1}\dots\beta_{n}$ are incident on a common vertex for every large enough $n\in\mathbb{N}$. In this case, we write $\alpha\sim\beta$.
This relation $\sim$ is supposed to glue together those sequences that represent the same point in the limit space. If two distinct sequences $\alpha$ and $\beta$ of $\Omega_{\mathcal{R}}$ are glued by $\sim$, then they must share a (possibly trivial) common prefix $x$, after which $x_{1}\dots x_{k}\alpha_{k+1}\dots\alpha_{m}$ and $x_{1}\dots x_{k}\beta_{k+1}\dots\beta_{m}$ must be incident on a common vertex for all $m\geq k$. For example, in the interval replacement system $\mathcal{I}$, for every finite prefix $p$ one has $p0\bar{1}\sim p1\bar{0}$, and these are all the gluings on $\mathcal{I}$.
@@
    there is no edge between $\iota_{c}$ and $\tau_{c}$ in any of the replacement graphs;
@@
    each replacement graph has at least two edges and at least one vertex that is not a boundary vertex;
This condition needs to be checked on the original replacement system, before applying the modification described in Subsection 1.1.1 that guarantees 1.2.
@@
By [BF19, Theorem 1.25], limit spaces are non-empty compact metrizable spaces. The limit space for the replacement system $\mathcal{I}$ from Figure 1 is the (topological) interval. The limit space of $\mathcal{D}_{n}$ from Figure 2 is the Ważewski dendrite $D_{n}$ of order $n$, and Figure 6(a)aaaimage in Figure 6(a) by Adam Majewski, licensed under CC BY 3.0 depicts $D_{3}$; see [dendrite] for more details on these spaces and their rearrangement groups. The limit space of $\mathcal{B}$ from Figure 4 (or, equivalently, Figure 3) is the Basilica limit space, which is homeomorphic to the Julia set for the quadratic polynomial map $z^{2}-1$, depicted in Figure 6(b)bbbimage in Figure 6(b) by Jim Belk and Bradley Forrest from [BF15]; see [BF15, BF19] for more details on the limit space and its rearrangement group.
@@ 2. The Gluing Automaton
Let $\Omega$ be a set of infinite sequences. We say that an equivalence relation on $\Omega$ is rational when there exists an automaton that recognizes a pair $(w,v)$ if and only if the two elements $w,v\in\Omega$ are related.
@@
Now consider an expanding replacement system $\mathcal{R}$ with colors $\mathrm{Col}=\{1,\dots,k\}$, base graph $\Gamma_{0}$ and replacement graphs $\Gamma_{1},\dots,\Gamma_{k}$. If the replacement system has loops, from here on we will always consider the replacement system where the boundary vertices for colors associated to loops are identified, as explained in Subsection 1.1.1. The gluing automaton $\mathrm{Gl}_{\mathcal{R}}=(\Sigma^{2},Q,\rightarrow,q_{0}(0))$ is built as follows.
@@
When processing one of the two entries of the $l$-th letter of a word in this alphabet, it will be natural to refer to that entry as if it were an edge of the $l$-th graph of the full expansion sequence (Definition 1.4). In this way, the two entries of the $l$-th letter of a word in $\Sigma^{2}$ represent two edges of the same graph, so in what follows we will simply describe their possible adjacency in the $l$-th graph of the full expansion sequence without explicitly mentioning the graph itself.
@@
The following Subsection will describe in more detail what these states mean. The intuition is that, while processing the edges of our two sequences, the first set $Q_{0}$ consists of an initial state $q_{0}(0)$ along with a state $q_{0}(i)$ for each color $i$ that indicates that the two entries being processed are the same $i$-colored edge, while the second set $Q_{1}$ has a state for each quadruple $\big{(}\begin{smallmatrix}i&\gamma\\ j&\delta\end{smallmatrix}\big{)}$ where $i,j$ are the colors of the edges and $\gamma,\delta$ distinguish between the types of adjacency between the edges. The symbols $\mathrm{lp}$ and $\mathrm{db}$ stand for loop and double, respectively, and their meaning will be made clear when describing the transition functions.
@@
$q_{0}(i)\xrightarrow{(a,b)}q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\alpha\\ \mathrm{col}(b)&\beta\end{smallmatrix}\big{)}$ for all $i\in\{0\}\cup\mathrm{Col}$ and $a\neq b$ in $\Gamma_{i}^{1}$, where $\alpha$ and $\beta$ depend on the adjacency between $a$ and $b$ in $\Gamma_{i}$ as shown in Table I.
@@
$q_{1}\big{(}\begin{smallmatrix}i&\gamma\\ j&\delta\end{smallmatrix}\big{)}\xrightarrow{(a,b)}q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\alpha\\ \mathrm{col}(b)&\beta\end{smallmatrix}\big{)}$ for all $i,j\in\{0\}\cup\mathrm{Col}$ whenever $a\in\Gamma_{i}^{1}$ and $b\in\Gamma_{j}^{1}$ satisfy the conditions described in Table II, which also depend on $\gamma$ and $\delta$ and determine $\alpha$ and $\beta$.
The meaning of these transitions (especially Type 2) will be made clear in Section 3. The idea is that Type 0 transitions simply induce a copy of the color graph (Definition 1.5), Type 1 transitions arise when two sequences first diverge and Type 2 transitions describe the adjacency of sequences after their divergence. Note that, when processing the pair $(a,b)$ in a state $q_{0}(i)$, there is a Type 1 transition if and only if $a$ and $b$ are adjacent in $\Gamma_{i}$.
Observe that loops at $\lambda_{i}$ may arise when the original replacement system (before gluing the initial and terminal vertices as explained in Subsection 1.1.1) had loops at $\iota$ or at $\tau$. Since the replacement system is expanding, this is the only possible occurrence of the column $\iota(a)=\tau(a)=\lambda_{i}$ (row $\iota(b)=\tau(b)=\lambda_{j}$).
Moreover, in Type 2 transitions $\alpha$ and $\beta$ will never be $\mathrm{db}^{+}$ nor $\mathrm{db}^{-}$, since the replacement system is expanding. Furthermore, note that when $\gamma$ (or $\delta$) is $\mathrm{db}^{+}$ or $\mathrm{db}^{-}$ one needs to check the incidence of $a$ (or $b$) with both the initial and terminal vertices $\iota_{i}$ and $\tau_{i}$ (or $\iota_{j}$ and $\tau_{j}$) of the replacement graph. However, since the replacement system is expanding, an edge of a replacement graph cannot be incident on both the initial and terminal vertices; thus, at most one of these can occur, so there is no ambiguity in the table.
$a$$b$ $a$$b$ $a$$b$ $a$$b$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{in}\\ \mathrm{col}(b)&\mathrm{out}\end{smallmatrix}\big{)}$ $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{in}\\ \mathrm{col}(b)&\mathrm{in}\end{smallmatrix}\big{)}$ $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{out}\\ \mathrm{col}(b)&\mathrm{out}\end{smallmatrix}\big{)}$ $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{out}\\ \mathrm{col}(b)&\mathrm{in}\end{smallmatrix}\big{)}$
$a$$b$ $a$$b$ $a$$b$ $a$$b$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{in}\\ \mathrm{col}(b)&\mathrm{lp}\end{smallmatrix}\big{)}$ $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{out}\\ \mathrm{col}(b)&\mathrm{lp}\end{smallmatrix}\big{)}$ $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{lp}\\ \mathrm{col}(b)&\mathrm{out}\end{smallmatrix}\big{)}$ $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{lp}\\ \mathrm{col}(b)&\mathrm{in}\end{smallmatrix}\big{)}$
$a$$b$ $a$$b$ $a$$b$
$q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{lp}\\ \mathrm{col}(b)&\mathrm{lp}\end{smallmatrix}\big{)}$ $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{db}^{+}\\ \mathrm{col}(b)&\mathrm{db}^{+}\end{smallmatrix}\big{)}$ $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\mathrm{db}^{+}\\ \mathrm{col}(b)&\mathrm{db}^{-}\end{smallmatrix}\big{)}$
Table I. Type 1 transitions and the types of adjacency.
@@
Table II. Type 2 transitions. Each entry $\begin{smallmatrix}\alpha\\ \beta\end{smallmatrix}$ represents a transition to the state $q_{1}\big{(}\begin{smallmatrix}\mathrm{col}(a)&\alpha\\ \mathrm{col}(b)&\beta\end{smallmatrix}\big{)}$. An empty entry at $\big{(}\begin{smallmatrix}\gamma\\ \delta\end{smallmatrix}\big{)}$ means that the state $q_{1}\big{(}\begin{smallmatrix}i&\gamma\\ j&\delta\end{smallmatrix}\big{)}$ does not exist.
For example, Figure 7 depicts the gluing automaton $\mathrm{Gl}_{\mathcal{I}}$ for the interval replacement system from Figure 1 (which is a classical example) and Figure 8 depicts the gluing automaton $\mathrm{Gl}_{\mathcal{D}_{n}}$ for the Dendrite replacement systems from Figure 2. Finally, Table III represents the gluing automaton of the Basilica replacement system: the automaton $\mathrm{Gl}_{\mathcal{B}}$ features a transition $q^{(1)}\xrightarrow{(a,b)}q^{(2)}$ if and only if Table III features a row that reads $q^{(1)}\mid(a,b)\mid q^{(2)}$.
For example, to deduce the transition $q_{1}\big{(}\begin{smallmatrix}1&\mathrm{in}\\ 1&\mathrm{out}\end{smallmatrix}\big{)}\xrightarrow{(1,0)}q_{1}\big{(}\begin{smallmatrix}1&\mathrm{in}\\ 1&\mathrm{out}\end{smallmatrix}\big{)}$ in Figure 7 from Table II one needs to observe that $\gamma=\mathrm{in}$ and $\tau(1)$ is the terminal vertex $\tau_{1}$ of the replacement graph, so one needs to consider the second column, and that $\delta=\mathrm{out}$ and $\iota(0)$ is the initial vertex $\iota_{1}$ of the replacement graph, so one needs to consider the fourth row. This entry of the table reads $\begin{smallmatrix}\mathrm{in}\\ \mathrm{out}\end{smallmatrix}$.
$q_{0}(0)$$q_{0}(1)$$q_{1}\big{(}\begin{smallmatrix}1&\mathrm{in}\\ 1&\mathrm{out}\end{smallmatrix}\big{)}$$q_{1}\big{(}\begin{smallmatrix}1&\mathrm{out}\\ 1&\mathrm{in}\end{smallmatrix}\big{)}$$(S,S)$$(0,0)$$(1,1)$$(0,1)$$(1,0)$$(1,0)$$(0,1)$
@@
$q_{0}(0)$$q_{0}(1)$$q_{1}\big{(}\begin{smallmatrix}1&\mathrm{out}\\ 1&\mathrm{out}\end{smallmatrix}\big{)}$$q_{1}\big{(}\begin{smallmatrix}1&\mathrm{in}\\ 1&\mathrm{in}\end{smallmatrix}\big{)}$$(S,S)$$(i,i)$$(i,j)$$(1,1)$$(n,n)$